The area of a rectangular wall of a barn is 260 square feet. Its length is 6 feet longer than twice its width. Find the length and width of the wall of the barn. THANK YOU!!
Answer by ikleyn(52809)

.
L*W = 260,

L = 2W + 6  

=================>

(2W+6)*W = 260

2W^2 + 6W - 260 = 0

W^2 + 3W - 130 = 0

(W+13)*(W-10) = 0


The only positive root makes sense:  W = 10.


Answer.  W = 10 ft;  L = 2W + 6 = 20+6 = 26 ft.